What is the internal rate of return per period for a project that requires an investment of 25,000 and returns 5,478 for 7 periods?
Answer #1

Calculating IRR from below formula

FV=PV *(1+R)^N

Given,

FV=5478,PV=25000,N=7

Putting the given information into the formula

5478=25000*(1+r)^7

5478/25000=(1+r)^7

0.2191=(1+r)^7

1+r = 0.805018

r = 0.195

IRR= 19.5%